Custom Cabinetry has one job in process (Job 120) as of June 30; at that time, its job cost sheet reports direct materials ofComplete this question by entering your answers in the tabs below. Required 1 Required 2 Prepare journal entries for the abovJournal entry worksheet 2 3 4 5 Record the entry to assign direct materials to job. Note: Enter debits before credits. Transa1 Record the entry to assign direct materials to job. 2 Record the entry to assign direct labor to job. 3 Record the applicatComplete this question by entering your answers in the tabs below. Required 1 Required 2 Compute the July 31 balances of the
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Custom Cabinetry
Calculation of Overhead cost applied in July Amount $ Note Note
Total labor cost in July                          10,500.00 A
Overhead rate 90% B
Overhead amount                            9,450.00 C=A*B
Calculation of cost of goods sold for Job 120
June July Total
Direct Materials                            7,500.00      2,700.00 10,200.00
Direct Labor                            3,300.00      2,300.00     5,600.00
Manufacturing Overhead                            2,970.00      2,070.00     5,040.00 Manufacturing Overhead for July is 90% of direct labor of July i.e. $ 2,300*90%= $ 2,070.
Cost of goods sold 20,840.00 D
Journal Entry
Transaction General Journal Debit $ Credit $
1 Work In Process 14,600.00
Direct Materials 14,600.00
(Being direct material used in production.)
2 Work In Process 10,500.00
Direct Labor 10,500.00
(Being direct labor used in production.)
3 Work In Process      9,450.00 See Note C
Manufacturing Overhead      9,450.00
(Being manufacturing overhead used in production.)
4 Accounts Receivable 23,500.00
Sales 23,500.00
(Being Job 120 sold on account.)
5 Cost of goods sold 20,840.00
Work In Process 20,840.00 See Note D
Calculation of Work In Process and Finished Goods
Work In Process Finished Goods
Job 122 Job 121
Direct Materials                            3,000.00      8,900.00
Direct Labor                            2,700.00      5,500.00
Manufacturing Overhead                            2,430.00      4,950.00 Manufacturing Overhead for July is 90% of direct labor of July i.e. $ 2,700*90% = $ 2,430 and $ 5,500*90%= $ 4,950.
Total                            8,130.00 19,350.00